# SLANG: New Concept Comprehension of Large Language Models

### 23 Jan 2024 v1

@@ Abstract
The dynamic nature of language, particularly evident in the realm of slang and memes on the Internet, poses serious challenges to the adaptability of large language models (LLMs). Traditionally anchored to static datasets, these models often struggle to keep up with the rapid linguistic evolution characteristic of online communities. This research addresses the critical need to bridge this gap, aiming to enhance LLMs’ comprehension of evolving new concepts on the internet, without the high cost and impracticality of continual retraining. To address this issue, we propose a new benchmark SLANG to assess LLMs’ proficiency in comprehending emerging linguistic trends and a baseline approach FOCUS, which uses causal inference to enhance LLMs to understand new phrases and usage patterns. This approach involves scrutinizing real-world instances of linguistic shifts, serving as contextual beacons, to form more precise and contextually relevant connections between newly emerging expressions and their intended meanings. The empirical analysis shows that our causal inference-based approach outperforms the traditional models in terms of precision and relevance in the interpretation of Internet slang and memes. [^1] Our code is available at https://github.com/Meirtz/FocusOnSlang-Toolbox.
@@ 1 Introduction
The evolution of human language is the history of human’s innovation and adaptation. From the earliest forms of communication to the complex linguistic systems of today, language has continuously transformed, reflecting, and shaping the human experience. Recently, the language evolution has been accelerated by the developing communication technology, particularly the Internet, which has introduced new dimensions to linguistic evolution (Varis and van Nuenen, 2017; Firth et al., 2019; Hammarström, 2016). These rapid changes in language poses serious challenges to the large language models (LLMs) on understanding the newly emerging concepts (Yang et al., 2023; Sun et al., 2021).
@@
Generally, LLMs are trained on static data (Brown et al., 2020), which limits their adaptivity to the dynamic and ever-evolving nature of human language. This limitation is particularly pronounced in the context of digital communication, where new forms of expression and knowledge emerge at an unprecedented pace (Sun et al., 2021). The crux of our research lies in exploring the potential of LLMs to adapt to and understand these linguistic transformations and new knowledge without resorting to continual retraining or fine-tuning with updated datasets or retrieving external data.
Note that many popular LLMs are often mainly black-box models. They may make decisions based on superficial patterns rather than justified reasons, which in turn can affect the model’s ability to understand and follow human instructions within context (Tang et al., 2023; Wang et al., 2022, Zhou et al., 2023b, Wang et al., 2023). For example, as shown in Figure 1, while CoT (Chain-of-Thought) Wei et al. (2022) provides a straightforward interpretation of the phrase "The sunset is beautiful, isn’t it?", our method recognizes the underlying metaphorical meaning, identifying it as an acknowledgment of an ending, such as a concluding phase in life or a relationship, in a nuanced conversational exchange. This recognition calls for a more nuanced approach in the development and evaluation of LLMs, which considers not only their performance metrics, but also the underlying principles guiding their decision-making processes.
In this study, we propose FOCUS (Factual cOntext CaUsal AnalysiS), a new methodology centered on causal inference, which can significantly enhance the language model’s capability to understand new phrases. This approach is grounded in the belief that understanding causal relationships in language usage, especially in evolving linguistic contexts, leads to more accurate and adaptable models. Our method involves constructing new datasets from UrbanDictionary Urban Dictionary LLC , comprising both factual and counterfactual instances. These datasets provide diverse linguistic contexts, ranging from common usage to generative scenarios, aiding in the comprehensive understanding of new phrases and concepts.
In addition to FOCUS, we introduce SLANG (Similarity of Lexical Analysis aNd Grasp), a new benchmark designed to evaluate language models’ adaptability to linguistic evolution and vocabulary changes. The SLANG benchmark focuses on assessing the capability of language models to maintain coherence and accuracy in the face of dynamic and unconventional language use, such as slang and idiomatic expressions, thereby evaluating our FOCUS methodology.
@@ 2 SLANG
In response to the rapid evolution of the Internet language, we introduce the SLANG benchmark, a tool derived from UrbanDictionary contributions. SLANG aims to evaluate the proficiency of LLMs in interpreting the ever-changing landscape of user-generated content. This benchmark is unique in its division into factual and counterfactual datasets, each serving a distinct purpose in understanding the adaptability of LLMs. The following subsections explore the definition of the SLANG problem, detail the dataset, and discuss the metrics used for evaluation. This structure ensures a comprehensive assessment of LLMs’ capabilities in navigating the complexities of modern linguistic trends.
2.1 Problem Definition
@@
2.2 Dataset Constructions
Data Extraction
@@ 2.2 Dataset Constructions
In the construction of the dataset, our methodology involved the systematic extraction of a substantial number of concepts from UrbanDictionary Urban Dictionary LLC , a widely recognized and reliable online repository of Internet slang and colloquial terms. UrbanDictionary stands out for its user-driven content, reflecting current language trends and the evolving nature of the internet lexicon. This platform serves as both a repository and a dynamic community, continuously enriching the evolving landscape of internet language with new concepts. Specifically, our approach involved selecting concepts that were added after a predetermined date, a strategy that was integral to ensuring the relevance and recency of the dataset. The subsequent steps undertaken for the development of this dataset were as follows. Our dataset’s construction commenced with UrbanDictionary as the foundational source. We meticulously extract relevant data, including the phrase, its user-provided definitions, usage examples, and user-generated ratings (upvotes and downvotes). This approach ensured that our dataset was grounded in authentic user-driven language trends. Additionally, the data construction pipeline is set up to automatically include fresh, non-member data for upcoming cutoff dates, ensuring the dataset stays up-to-date and comprehensive.
@@
Content Filtering
The content filtering of our dataset consists of several critical steps, with the foremost being the assurance of content novelty. Millions of concepts on UrbanDictionary have led to a significant proportion of phrases becoming commonplace in everyday communication. This prevalence implies that many of these phrases might already be included in GPT’s training corpus, as seen in the case of widely recognized terms like "ghosting" or "flex", which have transitioned from internet slang to general usage. As such, predicting the meanings of these phrases using LLMs would be redundant. Therefore, our primary objective was to ensure that the phrases and their definitions in our dataset were unfamiliar to the LLMs. In alignment with this objective, we factored in the official knowledge cut-off date of gpt-4-0613, which is April 2021 as detailed in the OpenAI documentation [^2] https://platform.openai.com/docs/models. To ensure the novelty and relevance of our dataset, we selected concepts that were added to UrbanDictionary after January 2022. This temporal gap was strategically chosen to include recent phrases that may have gained popularity online after the cut-off date but were not yet recorded in UrbanDictionary.
An essential aspect of content filtering is to analyze user-generated ratings, which we visualize in Figure 2. Comparative histograms in the figure illustrate the distribution of upvotes (left) and absolute upvotes (right) across dataset entries before and after our cleaning process. Both histograms are plotted on a logarithmic scale with the vertical axis representing the log percentage of the total dataset and the horizontal axis indicating the log count of upvotes or absolute upvotes. The blue bars represent the raw data, while the orange bars depict the cleaned data, facilitating a direct comparison of the distributions before and after data cleaning. This stringent data cleaning was crucial in refining the dataset to ensure the quality and reliability of the entries selected for our research.
We then rigorously employ gpt-4-0613 for an exhaustive filtration process, querying each term repeatedly to ensure that they were not memrozied by the GPT, thereby ensuring the unique and contemporary nature of our dataset.
Additionally, we rigorously filtered inappropriate content to uphold academic integrity, thoroughly examining each entry for explicit, violent, or harmful material to ensure the dataset’s appropriateness and quality.
Factual Dataset
Acknowledging the nonformal nature of the original dataset, we transformed the data into a uniform, formal dictionary format. This transformation involves simplifying the explanations and examples to make them universally understandable while preserving their original online meaning. The transformation process adhered to a custom-designed template, ensuring consistency and clarity across the concepts. However, considering that language models might output synonymous interpretations with varied wording, we augmented each transformed explanation with four additional variants incorporating synonym replacement. This approach was designed to capture the diversity of potential language model outputs. When evaluating predictions, we compare the model’s output with all five versions of the transformed explanations (the original plus four synonym-augmented variants) and select the result that most closely matches. This method ensures a more comprehensive and accurate assessment of the model’s capability to interpret and reproduce the factual content.
Counterfactual Dataset
Another core of our methodology lies in the creation of a counterfactual context. We aim to evaluate the adaptability and interpretative depth of LLMs in semantic decoding and contextual understanding. This approach is characterized by the preservation of original meanings from the factual dataset, while simultaneously introducing entirely new phrases, thereby shifting the linguistic form while maintaining the semantic substance.
This refined strategy requires the innovation of novel phrases that are linguistically distinct from their real-world counterparts, yet congruent with the retained meanings. The challenge lies in crafting these new phrases to be unique in structure, yet resonant with the established meanings, thus presenting a different verbal manifestation of the same conceptual core.
The process involves a creative and analytically rigorous approach, where the rich contextual foundation of the original dataset serves as the semantic basis for the generation of new phrases. These phrases are designed to encapsulate the same ideas as the original meanings, but articulated through a different linguistic prism. Achieving a balance between inventive language use and adherence to logical linguistic norms is crucial to ensuring that new phrases are plausible and coherent within established linguistic frameworks.
Upon generation, these new phrases are systematically paired with their corresponding original meanings, creating a dataset that juxtaposes novel linguistic expressions against established semantic structures. This pairing enables a multifaceted analysis of how LLMs process and comprehend language, specifically assessing their capability to disentangle linguistic form from semantic substance. This methodology offers insight into the nuanced capabilities of LLMs in interpreting language beyond mere recognition, exploring the realms of semantics and contextual linguistics.
@@ 2.3 Metrics
Given the nature of this task as a question-answering (QA) problem, traditional QA metrics like F1 score, Recall, and Precision were employed Yang et al. (2018). Moreover, we introduced BLEU (3-gram) Papineni et al. (2002) and ROUGE Lin (2004) as metrics to explore the capacity of the listed methods to generate higher quality interpretations with stricter similarity requirements. Again, considering that language models might output synonymous interpretations with varied wording, we also incorporated sentence similarity measures such as sentence-level similarity and SimCSE Gao et al. (2021), and sentence similarity Team (2023) is obtained by calculating the cosine similarity of the embeddings after the sentences are input into the model all-mpnet-base-v2 (Team, 2023; Song et al., 2020). A positive sample was considered for accuracy calculation if its SimCSE score exceeded 0.7. We also generated five lexically varied yet syntactically and semantically identical interpretations for each dataset entry. To determine the final metric, we selected the interpretation with the highest BLEU score.
@@ 3 FOCUS
Figure 3: Structural Causal Model (SCM) of large language models for interpreting new phrases, excluding confounders. The variables $X$ and $W$ encapsulate users’ complex intentions and thoughts, which span intricate emotional expressions, cultural insights, and extensive internet-specific knowledge. Grasping these nuanced aspects directly through an LLM is a challenging endeavor.
This section outlines the novel methodology employed in our research to enhance the adaptability of LLMs in understanding the evolving human language. Our method analogizes the dynamic nature of language to a continuously evolving entity that requires adaptive comprehension strategies.
@@ 3.1 Causal Analysis
3.1.1 Causal Analysis of Entity
Initially, we reference a typic SCM framework (Wang et al., 2023; Wang et al., 2022), assuming our SCM as $S=\{X,E,Y\}$, where $X$ denotes the context or input, and $E$ represents the confounder, encompassing phrases and other entities within the context. From the perspective of human understanding of new phrases, the interpretation $Y$ is derived from the context $X$ and its entities $E$. The confounder $E$ can be extracted from $X$, leading to the relationships $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$ in our model. By applying the $do$-operation Verma and Pearl (1990), which denoted as $do(X)$, we follow the guidelines (Wang et al., 2023; Wang et al., 2022) to conduct a rigorous causal inference, ensuring that the main effects of the textual context are captured without losing entity information. This operation aims to isolate the effect of the context $X$ on the confounder $E$. Consequently, the relationship between $X$ and $E$ undergoes a change, becoming $\tilde{X}\leftarrow\tilde{E}$, where $\tilde{E}$ represents the modified entity, and $\tilde{X}$ denotes the context obtained after the $do$ operation which effectively substitutes the actual entity.
Figure 5: The four-stage pipeline of FOCUS.
3.1.2 Analysis with Linguistic Factors
@@
Model Method F1 Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)

### 20 Feb 2024 v4

@@ Abstract
The dynamic nature of language, particularly evident in the realm of slang and memes on the Internet, poses serious challenges to the adaptability of large language models (LLMs). Traditionally anchored to static datasets, these models often struggle to keep up with the rapid linguistic evolution characteristic of online communities. This research aims to bridge this gap by enhancing LLMs’ comprehension of the evolving new concepts on the Internet, without the high cost of continual retraining. In pursuit of this goal, we introduce SLANG, a benchmark designed to autonomously integrate novel data and assess LLMs’ ability to comprehend emerging concepts, alongside FOCUS, an approach uses causal inference to enhance LLMs to understand new phrases and their colloquial context. Our benchmark and approach involves understanding real-world instances of linguistic shifts, serving as contextual beacons, to form more precise and contextually relevant connections between newly emerging expressions and their meanings. The empirical analysis shows that our causal inference-based approach outperforms the baseline methods in terms of precision and relevance in the comprehension of Internet slang and memes. [^1] Our code is available at https://github.com/Meirtz/FocusOnSlang-Toolbox.
@@ 1 Introduction
Recently, the language evolution has been accelerated by the online community, which has introduced new dimensions to linguistic shifts (Varis and van Nuenen, 2017; Firth et al., 2019; Hammarström, 2016). These rapid changes in language poses serious challenges to the large language models (LLMs) on understanding the newly emerging concepts (Yang et al., 2023; Sun et al., 2021).
Generally, LLMs are trained on static data (Brown et al., 2020), which limits their adaptivity to the dynamic and ever-evolving nature of human language. This limitation is particularly pronounced in the context of digital communication, where new forms of expression and concepts emerge at an unprecedented pace (Sun et al., 2021). Hence, it is essential for LLMs to to understand linguistic shifts and new concepts without constant updates or external data.
@@
Moreover, LLMs often make decisions based on superficial patterns rather than justified reasons. This can hinder their ability to accurately interpret and follow human instructions, as highlighted in several studies (Tang et al., 2023; Wang et al., 2022, Zhou et al., 2023b, Wang et al., 2023). For example, as depicted in Figure 1, the Chain-of-Thought (CoT) prompting Wei et al. (2022) simply interprets the phrase The sunset is beautiful, isn’t it? and misses the deeper, metaphorical meaning, which could represent an acknowledgment of an ending, like concluding a life phase or a relationship, in a complex conversation. This situation underscores the importance of enhancing and evaluating LLMs in a way that goes beyond their performance metrics. It’s essential to consider the fundamental principles that guide their decision-making processes.
Therefore, we propose SLANG (Similarity of Lexical Analysis aNd Grasp), a benchmark to assess language models’ adaptability to linguistic shifts, and FOCUS (Factual cOntext CaUsal analysiS), an approach based on causal inference for enhancing comprehension of new concepts.
The SLANG benchmark, developed from UrbanDictionary Urban Dictionary LLC , focuses on assessing the capability of language models to maintain coherence and accuracy in the face of dynamic and unconventional language use, such as slang and idiomatic expressions, thereby evaluating the capability of LLMs in grasping new concepts. We select recent entries after a specific cutoff date and filter out phrases already likely in LLM training data. We utilize user-generated ratings (upvotes and downvotes) to refine the dataset, ensuring its quality and comprehensiveness. The dataset is then standardized into a formal dictionary format, simplifying explanations and examples for universal understanding while retaining original meanings. This preprocessing approach ensures SLANG effectively evaluates LLMs’ adaptability to linguistic shifts.
FOCUS employs causal inference to enhance models’ comprehension of new concepts within evolving linguistic contexts. By analyzing causal relationships in language, FOCUS advances models’ predictive capabilities beyond traditional correlation-based learning. This method allows for a nuanced grasp of language dynamics, improving models’ adaptability and effectiveness in applications requiring deep understanding of language use. FOCUS significantly enhanced performance in language model comprehension, demonstrating superior precision and adaptability. With GPT-4, FOCUS achieved an $\text{F}_{1}$ score of 0.4532, precision of 0.4598, and recall of 0.4551, alongside an accuracy of 84.9%, outperforming previous methods in comprehension and adaptability.
The codes for the SLANG and FOCUS toolboxes are open-sourced, contributing to the community’s resources for advancing language model development.
@@ 2 SLANG
We introduce SLANG in response to rapidly evolving language. SLANG benchmark evaluates LLMs’ capability to interpret the dynamic landscape of user-generated new concepts. It uniquely features factual and counterfactual datasets, each crucial for gauging LLM adaptability. We detail SLANG’s problem definition, dataset construction, and evaluation metrics in the following subsections.
2.1 Problem
@@
2.2 Preprocessing
Extraction
@@ 2.2 Preprocessing
Our dataset construction involved extracting numerous concepts from UrbanDictionary, a platform known for user-generated content that reflects current language trends and the evolving internet lexicon, making it a unique, constantly updated repository and dynamic forum for new Internet language concepts. Specifically, our approach involved selecting concepts added after a predetermined date to ensure content novelty. We meticulously extract relevant data, including the phrase, its user-provided definitions, usage examples, and user-generated ratings (upvotes and downvotes). Additionally, the data construction pipeline is set up to automatically include fresh, non-member data for upcoming cutoff dates, ensuring the dataset stays up-to-date and comprehensive.
Filtering
The content filtering of our dataset consists of several steps, with the foremost being the assurance of content novelty. Many UrbanDictionary phrases are now common, potentially pre-included in GPT’s training data. Terms like ghosting or flex illustrate this shift to general usage. To avoid redundancy in LLM predictions, we focused on ensuring that the phrases and definitions in our dataset were new to the LLMs. To solve this, we factored in the official knowledge cut-off date of gpt-4-0613, which is April 2021 as detailed in the OpenAI documentation [^2] https://platform.openai.com/docs/models. To ensure the novelty and relevance of our dataset, we selected concepts that were added to UrbanDictionary after January 2022. This temporal gap was strategically chosen to include recent phrases that may have gained popularity online after the cut-off date but were not yet recorded in UrbanDictionary.
Another essential aspect of content filtering is to analyze user-generated ratings, which we visualize in Figure 2. Comparative histograms in the figure illustrate the distribution of upvotes (left) and absolute upvotes (right) across dataset entries before and after our cleaning process. This stringent data cleaning was crucial in refining the dataset to ensure the quality and reliability of the entries selected for our research.
Additionally, we employed gpt-4-0613 for thorough filtration, ensuring terms were not already memorized by GPT, thus maintaining our dataset’s novelty. We also removed inappropriate content, preserving academic integrity and ensuring quality.
Factual dataset
Acknowledging the informal nature of the original dataset, we transformed the data into a uniform, formal dictionary format. This transformation involves simplifying the explanations and examples to make them universally understandable while preserving their original online meaning. The transformation process adhered to a custom-designed template, ensuring consistency and clarity across the concepts. We also enriched each explanation with four synonym-based variants, ensuring the capture of diverse potential responses. When evaluating predictions, we compare the model’s output with all five versions of the transformed explanations (the original plus four synonym-augmented variants) and select the result that most closely matches. This method ensures a more comprehensive and accurate assessment of the model’s capability to interpret and reproduce the factual content.
Counterfactual dataset
We aim to assess the adaptability and interpretative depth of LLMs in semantic decoding and contextual understanding by preserving original meanings from the factual dataset while introducing new phrases. This approach maintains the semantic substance but shifts the linguistic form. Our strategy involves creating novel phrases that are linguistically distinct yet semantically congruent with their real-world counterparts. Crafting these phrases requires a creative and analytically rigorous approach, leveraging the rich contextual foundation of the original dataset. These new phrases encapsulate the same ideas as the original meanings but are expressed through a different linguistic perspective. Balancing inventive language use with logical linguistic norms is essential to ensure the new phrases are plausible and coherent. Upon generation, these new phrases are paired with their corresponding original meanings, creating a dataset that contrasts novel linguistic expressions against established semantic structures. This pairing allows for a multifaceted analysis of LLMs’ language processing and comprehension, particularly their ability to separate linguistic form from semantic substance. This methodology sheds light on the nuanced capabilities of LLMs in interpreting language beyond mere recognition, delving into semantics and contextual linguistics.
@@ 2.3 Metrics
For this task, we employed traditional metrics like $\text{F}_{1}$ score, recall, and precision Yang et al. (2018), and added BLEU (3-gram) Papineni et al. (2002) and ROUGE Lin (2004) for stricter quality checks. Considering that language models might output synonymous interpretations with varied wording, we also incorporated sentence similarity measures such as sentence-level similarity and SimCSE Gao et al. (2021), and sentence similarity Team (2023) as Similarity by calculating the cosine similarity of the embeddings after the sentences are input into the model all-mpnet-base-v2 (Team, 2023; Song et al., 2020). A positive sample was considered for accuracy calculation if its SimCSE score exceeded 0.7. We also generated five lexically varied yet syntactically and semantically identical interpretations for each dataset entry. To determine the final metric, we selected the interpretation with the highest BLEU score.
@@ 3 FOCUS
This section outlines the novel approach employed in our research to enhance the adaptability of LLMs in understanding the evolving human language. Our method analogizes the dynamic nature of language to a continuously evolving entity that requires adaptive comprehension strategies.
Figure 3: Structural Causal Model (SCM) of LLMs for interpreting new phrases, excluding confounders. The variables $X$ and $W$ encapsulate users’ complex intentions and thoughts, which span intricate emotional expressions, cultural insights, and extensive internet-specific knowledge. Grasping these nuanced aspects directly through an LLM is a challenging endeavor.
@@ 3.1 Causal Analysis
Analysis of entity
Initially, we reference a typical SCM framework (Wang et al., 2023; Wang et al., 2022), assuming our SCM as $S=\{X,E,Y\}$, where $X$ denotes the context or input, and $E$ represents the confounder, encompassing phrases and other entities within the context. From the perspective of human understanding of new phrases, the interpretation $Y$ is derived from the context $X$ and its entities $E$. The confounder $E$ can be extracted from $X$, leading to the relationships $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$ in our model. By applying the $do$-operation Verma and Pearl (1990), which denoted as $do(X)$, we follow the guidelines (Wang et al., 2023; Wang et al., 2022) to conduct a rigorous causal inference, ensuring that the main effects of the textual context are captured without losing entity information. This operation aims to isolate the effect of the context $X$ on the confounder $E$. Consequently, the relationship between $X$ and $E$ undergoes a change, becoming $\tilde{X}\leftarrow\tilde{E}$, where $\tilde{E}$ represents the modified entity, and $\tilde{X}$ denotes the context obtained after the $do$ operation which effectively substitutes the actual entity.
Analysis of linguistic factors
@@
3.2 Context-based Causal Intervention
Subsequently, we refine our causal intervention approach in the SCM, focusing on the role of context and linguistic factors. To concentrate on the contextual content and reduce reliance on shortcuts, $E$ is replaced with $\tilde{E}$, consequently transforming $X$ into $\tilde{X}$, (as illustrated in Figure 4 (a) and (b)). Further, to eliminate the entity bias in $W$, it is replaced with a specific placeholder, denoted as $\hat{W}$ (as illustrated in Figure 4 (c) and (d)). This alteration aids in isolating the effect of $W$ while excluding entity-specific biases. Overall, these adjustments accentuate the role of $R$ in shaping the context, enhancing the model’s capacity to highlight the influence of linguistic factors in a bias-free manner. The updated causal effect formula in the SCM is thus:
$P(Y=y|do(X=x))$
$P(\tilde{W}=\tilde{w}|do(X=x),\tilde{E}=\tilde{e},R=r)$
$P(\tilde{E}=\tilde{e})P(R=r)$
Figure 5: The four-stage pipeline of FOCUS.
The core idea of FOCUS is to explore how language models can adhere to guidelines to better understand the content of the context. The goal is to enable LLMs to analyze phrases according to usage examples and provide counterfactual interpretations, thereby understanding the evolving semantics of language. To achieve this objective, we propose a four-stage method (as shown in Figure 5), each with its principles:
Direct Inquiry (DI)
In the Direct Inquiry (DI) stage, we input the usage example (context) $X$ and phrase $W$ into the language model. Represented as $Y_{DI}=\mathcal{M}(X,W,E)$, this stage aims to evaluate the direct effect of $X$ on the inferred meaning $Y$. It allows the model to derive meaning naturally within the given context, focusing on the coherence of the phrase’s morphology, literal meaning, and context-based interpretation. DI sets a baseline for understanding the phrase’s meaning before adding more analytical layers.
Masked Entity Inquiry (MEI)
In the Masked Entity Inquiry (MEI) stage, we mask the phrase $W$ within context $X$ to analyze the meaning $Y$ without $W$’s direct influence. This process, $Y_{MEA},\hat{W}{MEA}=\mathcal{M}(X{\text{masked}},E)$, helps the model suggest synonyms or near-synonyms for the masked phrase. MEI focuses on extracting meaning from broader linguistic factors in $X$ and $E$, reducing bias towards $W$’s literal interpretation and enhancing context-based understanding.
Model Prompting Method $\textbf{F}_{\textbf{1}}$ Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)